# Reparameterized LLM Training via Orthogonal Equivalence Transformation

### v1

@@ 1 Introduction
  • Strong spectrum control: Because orthogonal transformations do not change the singular values of weight matrices, POET keeps the weight spectrum the same as the randomly initialized weight matrices (empirically validated by Figure 1 even with approximations). Through the initialization scheme, POET thus directly controls the singular value distribution of its weight matrices. As a result, and in contrast to standard LLM training, POET matrices avoid undesirable large singular values after training (Figure 1 and Appendix H). To further facilitate the POET algorithm, we introduce two new initialization schemes: normalized Gaussian initialization and uniform spectrum initialization, which can ensure the resulting weight matrices have bounded singular values.
@@
observe that the learning process can be divided into three distinct phases (Figure 2): (1) conical shell searching: The cosine starts at $1$ (i.e., $\bm{R}$ is the identity) and gradually converges to a stable range of $[0.6,0.65]$, which we observe consistently across all learnable orthogonal matrices. This suggests that $\bm{R}$ transforms $\bm{v}$ into a thin conical shell around its original direction. (2) stable learning on the conical shell: The cosine remains within this range while the model begins to learn stably. Despite the cosine plateauing, validation perplexity continues to improve almost linearly. (3) final adjusting: Learning slows and eventually halts as the learning rate approaches zero. We also find that training loss is generally not informative of these three phases. We provide an in-depth discussion and full empirical results in Appendix A,F. Our contributions are summarized below:
@@ 2 From Energy-preserving Training to Spectrum-preserving Training
Orthogonal training [46, 63, 49] is a framework to train neural networks by learning a layer-shared orthogonal transformation for neurons in each layer. Specifically, for a weight matrix $\bm{W}=\{\bm{w}_{1},\cdots,\bm{w}_{n}\}\in\mathbb{R}^{m\times n}$ where $\bm{w}_{i}\in\mathbb{R}^{m}$ is the $i$-th neuron, the layer’s forward pass is given by $\bm{y}=\bm{W}^{\top}\bm{x}$ with input $\bm{x}\in\mathbb{R}^{m}$ and output $\bm{y}\in\mathbb{R}^{n}$. Unlike standard training, which directly optimizes the weight matrix $\bm{W}$, orthogonal training keeps $\bm{W}$ fixed at its random initialization $\bm{W}_{0}={\bm{w}^{0}_{1},\ldots,\bm{w}^{0}_{n}}$ and instead learns an orthogonal matrix $\bm{R}\in\mathbb{R}^{m\times m}$ to jointly transform all neurons in the layer. The forward pass becomes $\bm{y}=(\bm{R}\bm{W}_{0})^{\top}\bm{x}$. The effective weight matrix in orthogonal training is $\bm{W}_{R}=\{\bm{w}^{R}_{1},\cdots,\bm{w}^{R}_{n}\}$ where $\bm{w}^{R}_{i}=\bm{R}\bm{w}_{i}$. A key property of orthogonal training is its preservation of hyperspherical energy. Letting $\hat{\bm{w}}_{i}=\frac{\bm{w}_{i}}{\|\bm{w}_{i}\|}$, orthogonal training ensures
@@
where hyperspherical energy $\text{HE}(\cdot)$ characterizes the hyperspherical uniformity of neurons by measuring the sum of pairwise similarities among them. Prior work [45, 72, 46, 47] has shown that energy-preserving training can effectively improve generalization. Orthogonal finetuning (OFT) [63, 49] further demonstrates that finetuning large foundation models while preserving hyperspherical energy achieves a favorable trade-off between efficient adaptation to downstream tasks and retention of pretraining knowledge. However, while the hyperspherical energy preservation is effective for finetuning, it can be overly restrictive for pretraining. To allow greater flexibility in the pretraining phase, we relax the constraint from preserving hyperspherical energy to preserving the singular-value spectrum instead. Because energy-preserving training inherently maintains the spectrum, it can be viewed as a special case of spectrum-preserving training. As a natural generalization, spectrum-preserving training learns a transformation $\mathcal{T}:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$ that perfectly preserves the spectrum, i.e.,
@@
Theorem 1 (Simplified informal results from [40]).
@@ 4 Discussions and Intriguing Insights
Guarantees of weight spectrum. For POET with standard and normalized Gaussian initializations, we prove in Appendix B that the largest and smallest singular values of weights can be bounded.
@@
The proof of this lemma can be found in Appendix C. The result extends to Block-stochastic SPO as this is strictly more expressive than fully stochastic SPO. The key idea of the proof is similar to the factorization of orthogonal matrices into a product of Givens rotations. Indeed, by multiplying $\bm{R}^{\top}$ with properly chosen primitive matrices $\bm{G}_{i}$ we can create zeros below the diagonal for one column after another. Note that each $\bm{G}_{i}$ has $b(b-1)/2$ parameters while $\bm{R}$ has $m(m-1)/2$ parameters, which implies that generally at least $\Omega((m/b)^{2})$ primitives are necessary. In Appendix C we also provide a heuristic that with high probability for $c=O(\ln(m)(m/b)^{2})$ every orthogonal matrix can be written as a product of $c$ orthogonal primitives $\bm{G}_{i}$.
@@ 5.1 LLM Pretraining using LLaMA Transformers
To highlight POET’s non-trivial performance improvement, we increase the training steps (i.e., effectively tokens seen) for AdamW, and find that POET-FS ($b$=1/2) still outperforms AdamW even even if AdamW is trained with almost triple the number of tokens. Results are given in Figure 7. In this experiment, the AdamW learning rate was carefully tuned for the full training run, and no training tokens were repeated. Thus, the improvement is non-trivial and cannot be attributed to merely increasing training steps. Interestingly, we also observe from Table 2 that POET’s performance appears strongly correlated with the parameter budget and larger budgets consistently yield better results across model scales. This is particularly important for model scaling law [33]. Another notable observation is that POET significantly outperforms LoRA [24] given a similar parameter budget. For instance, with approximately 30M trainable parameters, POET attains a validation perplexity of $18.05$, significantly better than LoRA’s $25.19$. We further observe that the block-stochastic variant is more parameter-efficient than the fully stochastic one. On the 130M model, it achieves a validation perplexity of $19.88$ with nearly 6M fewer trainable parameters, compared to $19.94$ for the fully stochastic variant. We hypothesize that this is due to better coverage of weight parameters. Specifically, the block-stochastic variant ensures all corresponding weights are updated at each step, unlike the more uneven updates in the fully stochastic variant. Experimental details and results on weight update coverage are provided in Appendix G.
@@ A.1 More Details on Vector Probing
The three training phases of POET are summarized from the empirical observation of the vector probing results. The idea of vector probing is very straightforward. We generate a constant vector $\bm{v}$ that is randomly initialized. Then we let it to be transformed by the learned orthogonal matrices $\bm{R}$ and $\bm{P}$. Finally, we compute the cosine of their angle: $\bm{v}^{\top}\bm{R}\bm{v}$ and $\bm{v}^{\top}\bm{P}\bm{v}$. In this process, the probing vector $\bm{v}$ is always fixed. The full results are given in Appendix F.
@@ A.3 Empirical Observations
Appendix B Guarantees of Weight Spectrum under POET
@@
Appendix C Proofs of Lemma 1
@@
Appendix D Experimental Details
@@
Appendix E Implementation and CUDA Acceleration
@@
Appendix F Results of Vector Probing for $\bm{R}$ and $\bm{P}$
@@
Appendix G Weight Update Evenness of Different POET Variants
@@
Appendix H Training Dynamics of Singular Values
@@
Appendix I Orthogonality Approximation Quality using Neumann Series
@@
Appendix J Full Results of Training Dynamics

### v2

@@ 1 Introduction
  • Strong spectrum control: Because orthogonal transformations do not change the singular values of weight matrices, POET keeps the weight spectrum the same as the randomly initialized weight matrices (empirically validated by Figure 1 even with approximations). Through the initialization scheme, POET thus directly controls the singular value distribution of its weight matrices. As a result, and in contrast to standard LLM training, POET matrices avoid undesirable large singular values after training (Figure 1 and Appendix I). To further facilitate the POET algorithm, we introduce two new initialization schemes: normalized Gaussian initialization and uniform spectrum initialization, which can ensure the resulting weight matrices have bounded singular values.
@@
observe that the learning process can be divided into three distinct phases (Figure 2): (1) conical shell searching: The cosine starts at $1$ (i.e., $\bm{R}$ is the identity) and gradually converges to a stable range of $[0.6,0.65]$, which we observe consistently across all learnable orthogonal matrices. This suggests that $\bm{R}$ transforms $\bm{v}$ into a thin conical shell around its original direction. (2) stable learning on the conical shell: The cosine remains within this range while the model begins to learn stably. Despite the cosine plateauing, validation perplexity continues to improve almost linearly. (3) final adjusting: Learning slows and eventually halts as the learning rate approaches zero. We provide an in-depth discussion and empirical results in Appendix A,G. Our major contributions are:
@@ 2 From Energy-preserving Training to Spectrum-preserving Training
Orthogonal training [47, 64, 50] is a framework to train neural networks by learning a layer-shared orthogonal transformation for neurons in each layer. For a weight matrix $\bm{W}=\{\bm{w}_{1},\cdots,\bm{w}_{n}\}\in\mathbb{R}^{m\times n}$ where $\bm{w}_{i}\in\mathbb{R}^{m}$ is the $i$-th neuron, the layer’s forward pass is $\bm{y}=\bm{W}^{\top}\bm{x}$ with input $\bm{x}\in\mathbb{R}^{m}$ and output $\bm{y}\in\mathbb{R}^{n}$. Unlike standard training, which directly optimizes $\bm{W}$, orthogonal training keeps $\bm{W}$ fixed at its random initialization $\bm{W}_{0}={\bm{w}^{0}_{1},\ldots,\bm{w}^{0}_{n}}$ and instead learns an orthogonal matrix $\bm{R}\in\mathbb{R}^{m\times m}$ to jointly transform all neurons in the layer. The forward pass becomes $\bm{y}=(\bm{R}\bm{W}_{0})^{\top}\bm{x}$. The effective weight matrix is $\bm{W}_{R}=\{\bm{w}^{R}_{1},\cdots,\bm{w}^{R}_{n}\}$ where $\bm{w}^{R}_{i}=\bm{R}\bm{w}_{i}$. A key property is its preservation of hyperspherical energy. With $\hat{\bm{w}}_{i}=\bm{w}_{i}/\|\bm{w}_{i}\|$, orthogonal training ensures
@@
where hyperspherical energy $\text{HE}(\cdot)$ characterizes the uniformity of neurons on a unit hypersphere. Prior work [46, 73, 47, 48] has shown that energy-preserving training can effectively improve generalization. Orthogonal finetuning (OFT) [64, 50] also demonstrates that finetuning foundation models while preserving hyperspherical energy achieves a favorable trade-off between efficient adaptation to downstream tasks and retention of pretraining knowledge. While the hyperspherical energy preservation is effective for finetuning, it can be too restrictive for pretraining. To allow greater flexibility in the pretraining phase, we relax the constraint from preserving hyperspherical energy to preserving the singular-value spectrum instead. By inherently maintaining the spectrum, energy-preserving training is a special case of spectrum-preserving training. As a generalization, spectrum-preserving training learns a transformation $\mathcal{T}:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$ that preserves the spectrum:
@@
Theorem 1 (Simplified results from [41]).
@@
POET preserves hyperspherical energy under isotropic Gaussian initialization. [47] shows that weight matrices that are initialized by zero-mean isotropic Gaussian distribution (e.g.[16]) are guaranteed to have small hyperspherical energy. Because zero-mean isotropic Gaussian is invariant under orthogonal transformation, $\bm{R}\bm{W}\bm{P}$ also has small hyperspherical energy (see Appendix B).
@@ 4 Discussions and Intriguing Insights
Guarantees of weight spectrum. For POET with standard and normalized Gaussian initializations, we prove in Appendix C that the largest and smallest singular values of weights can be bounded.
@@
The proof of this lemma can be found in Appendix D. The result extends to Block-stochastic SPO as this is strictly more expressive than fully stochastic SPO. The key idea of the proof is similar to the factorization of orthogonal matrices into a product of Givens rotations. Indeed, by multiplying $\bm{R}^{\top}$ with properly chosen primitive matrices $\bm{G}_{i}$ we can create zeros below the diagonal for one column after another. Note that each $\bm{G}_{i}$ has $b(b-1)/2$ parameters while $\bm{R}$ has $m(m-1)/2$ parameters, which implies that generally at least $\Omega((m/b)^{2})$ primitives are necessary. In Appendix D we also provide a heuristic that with high probability for $c=O(\ln(m)(m/b)^{2})$ every orthogonal matrix can be written as a product of $c$ orthogonal primitives $\bm{G}_{i}$.
@@ 5.1 LLM Pretraining using LLaMA Transformers
To highlight POET’s non-trivial performance improvement, we increase the training steps (i.e., effectively tokens seen) for AdamW, and find that POET-FS ($b$=1/2) still outperforms AdamW even even if AdamW is trained with almost triple the number of tokens. Results are given in Figure 7. In this experiment, the AdamW learning rate was carefully tuned for the full training run, and no training tokens were repeated. Thus, the improvement is non-trivial and cannot be attributed to merely increasing training steps. Interestingly, we also observe from Table 2 that POET’s performance appears strongly correlated with the parameter budget and larger budgets consistently yield better results across model scales. This is particularly important for model scaling law [34]. Another notable observation is that POET significantly outperforms LoRA [25] given a similar parameter budget. For instance, with approximately 30M trainable parameters, POET attains a validation perplexity of $18.05$, significantly better than LoRA’s $25.19$. We further observe that the block-stochastic variant is more parameter-efficient than the fully stochastic one. On the 130M model, it achieves a validation perplexity of $19.88$ with nearly 6M fewer trainable parameters, compared to $19.94$ for the fully stochastic variant. We hypothesize that this is due to better coverage of weight parameters. Specifically, the block-stochastic variant ensures all corresponding weights are updated at each step, unlike the more uneven updates in the fully stochastic variant. Experimental details and results on weight update coverage are provided in Appendix H.
@@ A.1 More Details on Vector Probing
The three training phases of POET are summarized from the empirical observation of the vector probing results. The idea of vector probing is very straightforward. We generate a constant vector $\bm{v}$ that is randomly initialized. Then we let it to be transformed by the learned orthogonal matrices $\bm{R}$ and $\bm{P}$. Finally, we compute the cosine of their angle: $\bm{v}^{\top}\bm{R}\bm{v}$ and $\bm{v}^{\top}\bm{P}\bm{v}$. In this process, the probing vector $\bm{v}$ is always fixed. The full results are given in Appendix G.
@@ A.3 Empirical Observations
Appendix B Minimum Hyperspherical Energy in POET
We start by showing that orthogonal equivalence transformation in POET can provably obtain small hyperspherical energy for its transformed weight matrices. This result holds when the weight matrices are independently initialized by zero-mean isotropic Gaussian distribution. Both Xavier [16] and Kaiming [22] initializations satisfy such a weight initialization condition. Orthogonal equivalence transformation is given by $\text{OET}(\bm{W};\bm{R},\bm{P})=\bm{R}\bm{W}\bm{P}$, where the input matrix $\bm{W}$ is multiplied on the left and on the right by orthogonal matrices $\bm{R}$ and $\bm{P}$, respectively.
When $\bm{W}$ is initialized by zero-mean isotropic Gaussian distribution, [47] has shown that these random neurons, if normalized, are uniformly distributed on the unit hypersphere. This leads to a provably small hyperspherical energy for the randomly initialized weight matrix. In the following, we will show that after orthogonal equivalence transformation, the weight matrix still maintains a small hyperspherical energy.
We consider a weight matrix $\bm{W}\in\mathbb{R}^{m\times n}$ where each entry is i.i.d. sampled from a zero-mean Gaussian distribution with variance the same variance $\sigma^{2}$, i.e., $W_{ij}\sim\mathcal{N}(0,\sigma^{2})$. After applying orthogonal equivalence transformation, we have $\bm{W}^{\text{new}}=\bm{R}\bm{W}\bm{P}$ where $\bm{R}$ and $\bm{P}$ are two orthogonal matrices. Then we compute the distribution of $\bm{W}^{\text{new}}$. Because linear maps preserve Gaussianity, each entry of $\bm{W}^{\text{new}}$ is a finite linear combination of $W_{ij}$, and hence, $\bm{W}^{\text{new}}$ follows a joint Gaussian which can be fully characterized by mean and covariance.
The mean of $\bm{W}^{\text{new}}$ is given by
$\mathbb{E}[\bm{W}^{\text{new}}]=\bm{R}\cdot\mathbb{E}[\bm{W}]\cdot\bm{P}=\bm{R}\cdot\bm{0}\cdot\bm{P}=\bm{0}=\mathbb{E}[\bm{W}].$ (15)
For its covariance, we consider two generic entries $W^{\text{new}}_{ij}$ and $W^{\text{new}}_{i^{\prime}j^{\prime}}$:
$W^{\text{new}}_{ij}$ $=\sum_{k,l}R_{ik}W^{\text{new}}_{kl}P_{lj},$ (16)
$W^{\text{new}}_{i^{\prime}j^{\prime}}$ $=\sum_{u,v}R_{i^{\prime}u}W^{\text{new}}_{uv}P_{vj^{\prime}}.$
Then we compute the covariance between the two entries:
$\text{Cov}(W^{\text{new}}_{ij},W^{\text{new}}_{i^{\prime}j^{\prime}})$ $=\sum_{k,l,u,v}R_{ik}R_{i^{\prime}u}P_{lj}P_{vj^{\prime}}\text{Cov}(W_{kl},W_{uv})$ (17)
$=\sigma^{2}(\bm{R}\bm{R}^{\top})_{ii^{\prime}}(\bm{P}^{\top}\bm{P})_{jj^{\prime}}$
$=\sigma^{2}\delta_{ii^{\prime}}\delta_{jj^{\prime}}$
which implies the following resutls:
  • The covariance matrix is a diagonal matrix, so different entries of $\bm{W}^{\text{new}}$ are uncorrelated.
  • Because $\bm{W}^{\text{new}}$ is a joint Gaussian and different entries are uncorrelated, each entry of $\bm{W}^{\text{new}}$ is independent.
  • Each entry of $\bm{W}^{\text{new}}$ has identical variance $\sigma^{2}$.
To sum up, each entry of $\bm{W}^{\text{new}}$ is i.i.d. $\mathcal{N}(0,\sigma^{2})$, which is identical to the distribution of each entry of $\bm{W}$. Because we have $\bm{W}^{\text{new}}=_{d}\bm{W}$, we can conclude that, similar to $\bm{W}$, $\bm{W}_{\text{new}}$ also has provably small hyperspherical energy among neurons.
Despite being extremely simple, we find that this is in fact a significant result. Under zero-mean isotropic Gaussian initialization, spectrum-preserving training and energy-preserving training can be achieved simultaneously. It also partially explains why the proposed normalized Gaussian initialization achieves the best performance.
Appendix C Guarantees of Weight Spectrum under POET
@@
Appendix D Proofs of Lemma 1
@@
Appendix E Experimental Details
@@
Appendix F Implementation and CUDA Acceleration
@@
Appendix G Results of Vector Probing for $\bm{R}$ and $\bm{P}$
@@
Appendix H Weight Update Evenness of Different POET Variants
@@
Appendix I Training Dynamics of Singular Values
@@
Appendix J Orthogonality Approximation Quality using Neumann Series
@@
Appendix K Full Results of Training Dynamics